OpenStudy (donnie1999):

HELLLLP ME PLEASE!!!!!! if a line with a slope of 4 crosses the y-axis at (0,1), what is the equation of the line?

OpenStudy (anonymous):

Well when we write out the equation of the line, we look for 2 things. The slope and the y-intercept. The y-intercept is what the y value will be when x is 0. Are you with me so far?

OpenStudy (donnie1999):

Yes I guess

OpenStudy (anonymous):

So we are going to write it out in the form \[\large y=mx+b\] with m as the slope and b as the y-intercept. What do you think it would look like in this form?

OpenStudy (donnie1999):

y=m*0+b

OpenStudy (anonymous):

Well m is the slope. So it HAS to be 4, right?

OpenStudy (donnie1999):

ok so slope is 4 but im still confused

OpenStudy (anonymous):

Here, let's take out the b for a second. \[\large y=4x\] So what happens when you plug in 0 for x?

OpenStudy (anonymous):

Na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, Batman! Na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, Batman!

OpenStudy (donnie1999):

The choices are y-4x=1 y+4x=1 x-4y=1 x+4y=1

OpenStudy (donnie1999):

so is it y+4X=1

OpenStudy (anonymous):

Na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, Batman! Na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, na, Batman!

OpenStudy (donnie1999):

@LegendarySadist can you help me out some more

OpenStudy (anonymous):

Yes, sorry for the delay. So if we have \[\large y=4x\] what will we get if we plug in 0 for x?

OpenStudy (anonymous):

you get babies

OpenStudy (donnie1999):

just x right because 4*0=0

OpenStudy (anonymous):

Yes so \[\large y=4*0\\\large y=0\] So when x is 0, y is 1. That means our y-intercept is 1. So now that we know that, how would we write the equation?

OpenStudy (donnie1999):

y-4x=1

OpenStudy (anonymous):

Yes, that would be it. Good job :)